Ali walks 120 meter east, 150 meter south 40 meter east find the distance and displacement?
1
Expert's answer
2020-02-17T09:16:46-0500

The distance


"d=d_1+d_2+d_3=120+150+40=310\:\rm m."

The displacement


"{\bf s}={\bf s_1}+{\bf s_2}+{\bf s_3}""=120{\bf i}+(-150){\bf j}+40{\bf i}=160{\bf i}-150{\bf j}."

The absolute value of displacement


"s=\sqrt{160^2+(-150)^2}=219\:\rm m."